OpenStudy (diyadiya):

Integrate \[\int \frac{x \ \cos^{-1} x}{\sqrt{(1-x^2)}}\]

OpenStudy (diyadiya):

By parts

OpenStudy (diyadiya):

I took Cos^-1x as 1st function & x/sqrt(1-x^2) as 2nd

myininaya (myininaya):

\[y=\cos^{-1}(x) => \cos(y)=x => -\sin(y) dy= dx\] Recall cos(y)=x/1 (=adj/hyp) pretend we drew a right triangle labeling the adj side to angle y x the hyp 1 and we use the Pythagorean thm to find the opposite side to angle y \[- \sqrt{1-x^2} dy =dx => dy=\frac{dx}{-\sqrt{1-x^2}}\] So we have \[-\int\limits_{}^{}\cos(y) \cdot y dy\] I like this form to integrate more...

myininaya (myininaya):

\[-\sin(y) \cdot y - \int\limits_{}^{} -\sin(y) dy=-\sin(y) \cdot y-\cos(y)+C\]

OpenStudy (diyadiya):

Hm okay But the answer is different

myininaya (myininaya):

\[- \frac{\sqrt{1-x^2}}{1} \cdot \cos^{-1}(x)-x+C\]

myininaya (myininaya):

Using the right triangle I was able to put it back in terms of x

OpenStudy (diyadiya):

Lol yeah ,Thats the answer

OpenStudy (diyadiya):

Can you show me another method ? If you dont mind , this is confusing me a bit

OpenStudy (diyadiya):

Its okay wait , i'll just go through this again

OpenStudy (diyadiya):

\[- \sqrt{1-x^2} dy =dx \] How did you get this ?

myininaya (myininaya):

The part where I drew a right triangle?

myininaya (myininaya):